# Prompt2DEM: High-Resolution DEMs for Urban and Open Environments from Global Prompts Using a Monocular Foundation Model

### v1

@@ 1 Introduction
In response to the pressing need for high-resolution DEMs (HR-DEMs), super-resolution (SR) approaches have been developed and refined to enhance the spatial resolution of existing DEMs using machine learning (ML) or spatial interpolation techniques Zhang and Yu (2022a). As a next step, recently developed convolutional neural networks (CNNs), generative adversarial networks (GANs), and vision transformers (ViTs) were shown to be able to generate HR-DEMs from low-resolution data sources, significantly outperforming previous spatial ML and interpolation methods (Wang et al., 2024b). More recent studies have explored the integration of auxiliary modal information from remote sensing imagery for deep learning DEM super-resolution, such as shaded relief (Huang et al., 2024), hydrological knowledge (Cao et al., 2025) and relative depth (Huang et al., 2025).
@@
    We enhance the spatial resolution of predicted DEMs by a factor of 100, from 30 m to 30 cm, surpassing previous attempts by an order of magnitude and achieving up to an 18 % improvement in vertical accuracy compared with the original SRTM dataset.
@@
2 Related Work
This section reviews recent advances and existing research gaps in methods aimed at improving the spatial resolution of DEMs. It focuses on three main approaches: super-resolution techniques applied to DEMs (DEMSR); guided DEMSR methods that leverage deep learning to refine low-resolution DEMs using auxiliary data; and the application of monocular depth estimation to improve DEMs.
2.1 DEMSR
Early attempts to apply deep learning SR to DEMs was based on CNNs that were trained to produce high-resolution DEMs (HR-DEMs) from low-resolution DEMs (LR-DEMs) based on learned relations Demiray et al. (2021). For example, D-SRCNN (Demiray et al., 2020) based on super-resolution convolutional neural networks (SRCNNs) (Dong et al., 2015), MSRResNet (Jiang et al., 2023) based on super-resolution residual convolutional neural networks (SRResNets) (Ledig et al., 2017) and TfaSR integrating terrain feature-aware super resolution with residual CNNs. Compared with the results of traditional interpolation methods such as bicubic interpolation and existing neural networks, the results obtained with TfaSR were improved by 7.8% (for an overall improvement of 53.7%) when the DEM spatial resolution was increased from 30 m to 10 m (Zhang et al., 2022).
Recently, more advanced enhanced super-resolution generative adversarial networks (ESRGANs) (Wang et al., 2018) were applied to recover a 30-m HR DEM from a 90-m LR DEM (Wu and Ma, 2020). Furthermore, ViTs have been proposed to function well in complex terrains Li et al. (2023). For example, a ViT-based topography neural network (TTSR) was validated using two publicly available real-world DEM datasets to recover a DEM from 30 to 10 m. Also, the root mean square error (RMSE) for elevation accuracy was reduced by approximately 6%–30% compared to state-of the-art SR methods (Wang et al., 2024b).
In line with the ongoing development of DEMSR, studies in the past two years have investigated the integration of auxiliary modal information (Huang et al., 2024, 2025; Cao et al., 2025; Paul and Gupta, 2024). For example, the attention-based hierarchical terrain fusion (AHTF) framework uses shaded relief as its guiding modality and outperforms state-of-the-art methods (Gao et al., 2025). In another approach, hydrological knowledge derived from surface flow direction computations was integrated into a deep learning-based DEM super-resolution framework. The proposed method outperformed previous methods in generating 5 m resolution DEMs from 20-m resolution inputs (Cao et al., 2025).
2.2 Monocular Depth Estimation
Considering the aforementioned shortcomings of the different DEMSRs, deep learning-based MDE models allow depth to be quantified from a single RGB image using training sets (Zhao et al., 2020). Instead of using LR-to-HR translation to estimate depth, MDE models rely on scene semantics using RGB information (Yang et al., 2024a). Furthermore, MDE models are not constrained by fixed upscaling factors; rather, they produce depth estimates at the resolution of the RGB input.
To minimize loss during training, most MDEs use encoder–decoder architecture (Kim et al., 2018). Depth estimation relies on cues such as shading, occlusion, perspective, texture variations, and object scaling to differentiate and understand the scene (Godard et al., 2017). MDE has been used for 3D reconstruction, augmented reality, autonomous driving, and robotics (Ming et al., 2021). These uses of MDEs typically involve training a neural network model to learn the relationship between image pixels and their corresponding depth values.
Foundation models have recently gained traction in computer vision tasks (Lüddecke and Ecker, 2022), where they offer enhanced generalization by learning high-level concepts and relationships rather than being restricted to predefined class labels (Liu et al., 2024). Depth Anything V2, which is designed to be a foundation model for MDE, provides a highly practical solution for robust MDE. It is benchmarked on standard datasets, e.g., KITTI, Sintel, and more, and it outperforms previous foundation MDE models, such as MiDas V3.1 from Intel Labs (Birkl et al., 2023b). The superior performance of Depth Anything V2 was evident when evaluated on the DA-2K dataset, which includes eight different image categories, 9% of which are aerial imagery and fine-detail images (Yang et al., 2024b). These results indicate that terrain elevation could potentially be estimated from drone imagery.
To address the limitations related to global elevation recovery, PromptDA—drawing inspiration from a new paradigm introduced in its framework—integrates sparse depth information at multiple scales within the depth decoder. Prompt fusion has demonstrated that prompts can convey valuable metric information. As shown in recent evaluations, PromptDA outperforms all existing monocular depth estimation (MDE) methods (Lin et al., 2025).
@@ 3.2 Training and prompt engineering
Current depth foundation models (Yang et al., 2024a; Birkl et al., 2023a) generally share similar network structures to dense prediction transformer (DPT) (Ranftl et al., 2021) networks. Specifically, they use a vision transformer (ViT) with multiple stages to extract tokenized image features that represent the feature map at each stage. The DPT decoder reassembles features from different stages into image-like representations using the reassemble operation. Finally, a sequence of convolutional blending steps is applied to merge features across different stages and produce a dense depth map.
PromptDA proposes a concise prompt-fusion architecture that is tailored for a DINOv2 encoder, a DPT-based decoder, and a prompt injection mechanism to enable the integration of low-resolution (LR) depth information (Lin et al., 2025). Prompts are bilinearly resized to match the current scale of the spatial dimensions $R\in\mathbb{R}^{1\times H_{i}\times W_{i}}$. The resized depth map is then passed through a shallow convolutional network to extract depth features. These are projected to match the channel dimensions of image features $F_{i}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ using a zero-initialized convolutional layer. Subsequently, depth features are added to intermediate features of the DPT decoder for depth prediction. A sample of training patches is shown in Fig. 4.
Typically, training a prompt-based monocular model requires LR depth information as prompts and precise ground truth (GT) depth for backpropagation. Here, because the data are transitioned to elevation values, we trained the model using (1) the global DEM 30-m SRTM, and (2) HR LiDAR-derived DEMs (50 cm and 1 m). For accurate predictions, we unfreezed all weights during the initial training. This step was needed because the pretrained weights of the PromptDA model were trained on ground photography, captured by iPhone devices, and prompted with low-resolution depth data from the iPhone’s LiDAR sensor. Moreover, the prompts we used are fundamentally different because surface elevation data use positive values, i.e., above mean sea level. In contrast, MDE is typically used to estimate camera-object distances.
@@ 4 Results
The performance of the prompt-based framework was evaluated by comparing absolute terrain feature errors between predicted DEMs, SRTM, and HR-LiDAR reference data. The results show that the predicted DEM maintained the spatial resolution of the input RGB imagery (0.6 m), which provided substantially greater amounts of more valuable detail than was provided by the SRTM data (Figures 6, 7, and 8).
@@ 4.1 HR-DEM estimation
The metrics compared indicate that, in general, the prompt-based framework significantly improved the absolute accuracy of the DEM, and this effect was more noticeable in the vegetated and urban landscapes, although it also performed well in the simpler, bare soil landscapes (Figures 6A, 7A and 8A).
@@ 4.2 HR-DEM void filling
A comparative analysis of the different types of landscape showed that the accuracy of the elevation estimates was optimal in the urban and bare soil landscapes. In the vegetated landscapes, however, the elevation estimation was slightly less accurate relative to the DEM estimation benchmark. The improved accuracies of the slope and aspect estimations in vegetated areas paralleled those observed in the primary DEM estimation task. See Figures 6B, 7B, and 8B.
@@ 4.4 Seamless DEM
Stream network accuracy metrics for natural vegetated landscapes are presented in Table 6, while the corresponding metrics for bare landscapes are shown in Table 7. Seamless DEM reconstruction was achieved by unpatchifying predicted patches using a distance-based linear blending method. The resulting seamless DEMs for both sites are visualized in Figures 10 and 11. Stream networks were calculated in ArcGIS Pro and were compared with channel networks extracted from both SRTM and HR-LiDAR DEMs. Relative to the LiDAR data, the performance of the predicted DEM was consistently superior to that of the SRTM baseline data in almost all types of landscape.
@@ 5.1 Performance using different prompts
In this study, we observed that the source of the prompts affects prediction accuracy. Predictions improved as prompt RMSE decreased and spatial resolution increased. SRTM prompts led to less accurate predictions compared with fused prompts that used holed HR-DEMs combined with SRTM (see Table 3 vs Table 4). HR-DEMs provided more accurate absolute and local feature information, even outside the holed regions, resulting in improved predictions in the predicted voids, particularly in urban and bare landscapes. The lower performance obtained in vegetated landscapes can be attributed to the large discrepancy between the SRTM and the HR-DEM in the holed areas, which caused significantly discontinuous prompts. Moreover, the best predictions were made using HR-DTM prompts, which allowed the framework to estimate only the fine features, such as buildings and vegetation, while the terrain was provided with high accuracy (see Fig.  9 and Table 5).
@@ 5.3 MDE vs SR
Super-resolution (SR) approaches were used in a variety of studies to improve DEM (Zhang and Yu, 2022a). Although monocular depth estimation is attracting increased attention in remote sensing studies, they are still rarely used in studies of open environments, particularly in the context of hydrological and geomorphological processes such as hillslope erosion and stream network extraction (Tangellapalli et al., 2024; Rafaeli et al., 2025). Note that there is a critical conceptual difference between SR and MDE. Although MDE is superior to SR in terms of accuracy and resolution, it has been less widely adopted for DEM enhancement due to its lack of global context (Florea and Nedevschi, 2025), as discussed earlier in the Related Work section (Tang et al., 2021).
Recent attempts have been made to integrate MDE into super-resolution models. Depth Anything V2 incorporates relative depth, which adds valuable information to super-resolution prediction networks and achieves state-of-the-art accuracy. While this integration represents a meaningful step toward combining the two approaches, the foundation of the proposed model remains rooted in a super-resolution architecture rather than in MDE Huang et al. (2025).
However, SR methods have inherent limitations: (1) their design is inflexible because the training resolution is fixed. SR models are trained to map from a specific low-resolution (LR) input to a high-resolution (HR) output (Wang et al., 2018, 2024b); and (2) their upscaling operators are fixed because SR architectures are typically constrained to use convolutions or learned interpolations tailored to specific upscaling factors, usually up to 16× (Lepcha et al., 2023).
@@ 5.4 Large-scale application
HR-DEMs in the United States are provided primarily by the USDA Natural Resources Conservation Service (NRCS) and the USGS 3D Elevation Program (3DEP) (Survey, 2025). Nearly 80% of the contiguous United States (CONUS) has been mapped with 1-m LiDAR-derived DEMs , while for some regions, sub-meter DEMs are available, see Figure 16. In addition to the issue of spatial coverage, these HR-DEMs have temporal limitations. For example, annual nationwide updates have not been implemented, in contrast to the high frequency of RGB image acquisitions provided by the NAIP or by HR-satellite imagery (Sugarbaker et al., 2017).
Figure 16: HR-DEM availability according to the USGS 3D Elevation Program (3DEP), indicating where datasets are available at standardized horizontal resolutions. Source: nationalmap.gov/arcgis/rest/services/3DEPElevationIndex/MapServer/

### v2

@@ 1 Introduction
In response to the pressing need for high-resolution DEMs (HR-DEMs), super-resolution (SR) approaches have been developed and refined to enhance the spatial resolution of existing DEMs using machine learning (ML) or spatial interpolation techniques Zhang and Yu (2022a). As a next step, super-resolution convolutional neural networks (SRCNNs), super-resolution generative adversarial networks (ESRGANs), and super-resolution vision transformers (TTSR) were shown to be able to generate HR-DEMs from low-resolution data sources, significantly outperforming previous spatial interpolation methods (Wang et al., 2024b). More recent studies have explored the integration of auxiliary modal information from remote sensing imagery for deep learning DEM super-resolution, such as shaded relief (Huang et al., 2024), hydrological knowledge (Cao et al., 2025) and relative depth (Huang et al., 2025).
@@
    We introduce an elevation-prompted monocular foundation model for the first time.
@@ 2.1 Study area and dataset
2.2 Model architecture
Current depth foundation models (Yang et al., 2024a; Birkl et al., 2023a) generally share similar network structures to dense prediction transformer (DPT) (Ranftl et al., 2021) networks. Specifically, they use a vision transformer (ViT) with multiple stages to extract tokenized image features that represent the feature map at each stage.
ViT Encoder — In the proposed framework, we used a pretrained ViT encoder from the Depth Anything V2 model, which builds on the largest DINOv2 and leverages its inherent generalization ability (Yang et al., 2024c). In particular, Depth Anything V2 introduces several key training advantages: (1) It overcomes the limitations of manual image annotation by incorporating synthetically annotated data during training, with fine-detailed labels; (2) It employs a teacher–student architecture to obtain a lightweight yet accurate model. The model is initially trained on synthetic images, after which the teacher model generates pseudo masks on unannotated real images. Finally, a smaller student model is trained on these pseudo-annotated images; (3) Its annotation pipeline integrates SAM for image sampling, utilizing a combination of real and pseudo-labeled depth maps derived from 62 million images (Yang et al., 2024b).
Depth Anything V2 is benchmarked on standard datasets such as KITTI, Sintel, and others, and it outperforms previous foundation MDE models, including MiDaS v3.1 from Intel Labs (Birkl et al., 2023b). Its superior performance is further demonstrated on the DA-2K dataset, which includes eight different image categories, 9% of which consist of aerial and fine-detail imagery
DPT Decoder — The DPT decoder reassembles features from different stages into image-like representations using the reassemble operation. Finally, a sequence of convolutional blending steps is applied to merge features across different stages and produce a dense depth map, to produce full resolution predictions using a convolutional decoder. For monocular depth estimation, we observe an improvement of up to 28% in relative performance when compared to a state-of-the-art fully-convolutional network.Ranftl et al. (2021).
Prompt Fusion Block — PromptDA proposes a concise prompt-fusion architecture that is tailored for a DINOv2 encoder, a DPT-based decoder, and a prompt injection mechanism to enable the integration of low-resolution (LR) depth information (Lin et al., 2025). Prompts are bilinearly resized to match the current scale of the spatial dimensions $R\in\mathbb{R}^{1\times H_{i}\times W_{i}}$. The resized depth map is then passed through a shallow convolutional network to extract depth features. These are projected to match the channel dimensions of image features $F_{i}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ using a zero-initialized convolutional layer. Subsequently, depth features are added to intermediate features of the DPT decoder for depth prediction, see Fig 6.
Figure 6: Architecture of the Prompt Fusion Block. The low-resolution DEM is linearly interpolated, passed through a shallow convolutional network, and projected via a zero-initialized layer. The result is added to the intermediate DPT features to inject elevation context into the dense prediction transformer.
@@ 2.3 Training and prompt engineering
Typically, training a prompt-based monocular model requires LR depth information as prompts and precise ground truth (GT) depth for backpropagation. Here, because the data are transitioned to elevation values, we trained the model using (1) the global DEM 30-m SRTM, and (2) HR LiDAR-derived DEMs (50 cm and 1 m). For accurate predictions, we unfreezed all weights during the initial training. This step was needed because the pretrained weights of the PromptDA model were trained on ground photography, captured by iPhone devices, and prompted with low-resolution depth data from the iPhone’s LiDAR sensor. Moreover, the prompts we used are fundamentally different because surface elevation data use positive values, i.e., above mean sea level. In contrast, MDE is typically used to estimate camera-object distances. A sample of training patches is shown in Fig. 4.
@@ 3 Results
The performance of the prompt-based framework was evaluated by comparing absolute terrain feature errors between predicted DEMs, SRTM, and HR-LiDAR reference data. The results show that the predicted DEM maintained the spatial resolution of the input RGB imagery (0.6 m), which provided substantially greater amounts of more valuable detail than was provided by the SRTM data (Figures 7, 8, and 9).
@@ 3.1 HR-DEM estimation
The metrics compared indicate that, in general, the prompt-based framework significantly improved the absolute accuracy of the DEM, and this effect was more noticeable in the vegetated and urban landscapes, although it also performed well in the simpler, bare soil landscapes (Figures 7A, 8A and 9A).
@@ 3.2 HR-DEM void filling
A comparative analysis of the different types of landscape showed that the accuracy of the elevation estimates was optimal in the urban and bare soil landscapes. In the vegetated landscapes, however, the elevation estimation was slightly less accurate relative to the DEM estimation benchmark. The improved accuracies of the slope and aspect estimations in vegetated areas paralleled those observed in the primary DEM estimation task. See Figures 7B, 8B, and 9B.
@@ 3.4 Seamless DEM
Stream network accuracy metrics for natural vegetated landscapes are presented in Table 6, while the corresponding metrics for bare landscapes are shown in Table 7. Seamless DEM reconstruction was achieved by unpatchifying predicted patches using a distance-based linear blending method. The resulting seamless DEMs for both sites are visualized in Figures 11 and 12. Stream networks were calculated in ArcGIS Pro and were compared with channel networks extracted from both SRTM and HR-LiDAR DEMs. Relative to the LiDAR data, the performance of the predicted DEM was consistently superior to that of the SRTM baseline data in almost all types of landscape.
@@ 4.1 Performance using different prompts
In this study, we observed that the source of the prompts affects prediction accuracy. Predictions improved as prompt RMSE decreased and spatial resolution increased. SRTM prompts led to less accurate predictions compared with fused prompts that used holed HR-DEMs combined with SRTM (see Table 3 vs Table 4). HR-DEMs provided more accurate absolute and local feature information, even outside the holed regions, resulting in improved predictions in the predicted voids, particularly in urban and bare landscapes. The lower performance obtained in vegetated landscapes can be attributed to the large discrepancy between the SRTM and the HR-DEM in the holed areas, which caused significantly discontinuous prompts. Moreover, the best predictions were made using HR-DTM prompts, which allowed the framework to estimate only the fine features, such as buildings and vegetation, while the terrain was provided with high accuracy (see Fig.  10 and Table 5).
@@ 4.3 MDE vs SR
Super-resolution (SR) approaches were used in a variety of studies to improve DEM (Zhang and Yu, 2022a). Although monocular depth estimation is attracting increased attention in remote sensing studies, they are still rarely used in studies of open environments, particularly in the context of hydrological and geomorphological processes such as hillslope erosion and stream network extraction (Tangellapalli et al., 2024; Rafaeli et al., 2025). Note that there is a critical conceptual difference between SR and MDE. Although MDE is superior to SR in terms of accuracy and resolution, it has been less widely adopted for DEM enhancement due to its lack of global context (Florea and Nedevschi, 2025).
@@ 4.4 Large-scale application
HR-DEMs in the United States are provided primarily by the USDA Natural Resources Conservation Service (NRCS) and the USGS 3D Elevation Program (3DEP) (Survey, 2025). Nearly 80% of the contiguous United States (CONUS) has been mapped with 1-m LiDAR-derived DEMs , while for some regions, sub-meter DEMs are available (Survey, 2025). In addition to the issue of spatial coverage, these HR-DEMs have temporal limitations. For example, annual nationwide updates have not been implemented, in contrast to the high frequency of RGB image acquisitions provided by the NAIP or by HR-satellite imagery (Sugarbaker et al., 2017).